# Spin Chern number in altermagnets

### v1

@@ III.3 Bulk invariants for $C_{4}^{z}\mathbb{T}$ on 3D torus
${\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(U\cap V)\cong{{K}}^{0}_{C_{2}}(\{k_{z}=\tfrac{\pi}{2}\})$ (99)
@@
${\mathcal{K}}^{-1}_{C_{4}\mathbb{T}}(U\cap V)\cong{{K}}^{-1}_{C_{2}}(\{k_{z}=\tfrac{\pi}{2}\})=0,$ (101)
@@
$0\to{\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{3})\to{\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(U)\oplus{\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(V)\to{\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(U\cap V)$ (102)
@@
$0\to{\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{3})\to(\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2})\oplus(\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2})\to\mathbb{Z}^{\oplus 2}.$ (103)
@@
${\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{3})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ (104)
where the two $\mathbb{Z}_{2}$ components measure the $\mathbb{Z}_{2}$ invariant on each of the planes $k_{z}=0$ and $k_{z}=\pi$. Whenever the two invariants differ, which means that the spin Chern number changes between the two planes, then we know that the Chern-Simons axion insulator $\theta$ term is not zero 34.
@@
${\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{3})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ (105)
@@ IV Conclusions
Furthermore, we have extended this Hamiltonian to a 3D altermagnetic one preserving the $C_{4}^{z}\mathbb{T}$ symmetry. In the negligible SOC limit, our analysis reveals the presence of two Weyl points. However, when incorporating SOC, we observe the emergence of eight Weyl points and two spin Weyl points. In these cases, our model defines a Weyl altermagnetic semimetal, in which the spin Chern signal is linked to the locations of both the Weyl and spin Weyl points. The presence of these Weyl points confirms that the material exhibits non-trivial topological features characteristic of a Weyl semimetal, opening pathways to explore exotic phenomena such as the chiral anomaly and surface state physics.
@@ Acknowledgments
RGH gratefully acknowledges the computing time granted on the supercomputer Mogon at Johannes Gutenberg University Mainz (hpc.uni-mainz.de). HS acknowledges the support of CONAHCyT through grant number CVU 926934. BU acknowledges the support of Max Planck Institute for Mathematics in Bonn, Germany. RGH and BU thank the continuous support of the Alexander Von Humboldt Foundation, Germany.

### v2

@@ III.3 Bulk invariants for $C_{4}^{z}\mathbb{T}$ on 3D torus
${\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(U\cap V)\cong{{K}}^{0}_{C_{2}^{z}}(\{k_{z}=\tfrac{\pi}{2}\})$ (99)
@@
${\mathcal{K}}^{-1}_{C_{4}^{z}\mathbb{T}}(U\cap V)\cong{{K}}^{-1}_{C_{2}^{z}}(\{k_{z}=\tfrac{\pi}{2}\})=0,$ (101)
@@
$0\to{\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{3})\to{\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(U)\oplus{\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(V)\to{\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(U\cap V)$ (102)
@@
$0\to{\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{3})\to(\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2})\oplus(\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2})\to\mathbb{Z}^{\oplus 2}.$ (103)
@@
${\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{3})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ (104)
where the two $\mathbb{Z}_{2}$ components measure the $\mathbb{Z}_{2}$ invariant on each of the planes $k_{z}=0$ and $k_{z}=\pi$. Whenever the two invariants differ, which means that the spin Chern number changes between the two planes, then we know that the Chern-Simons axion insulator $\theta$ term is not zero 43.
@@
${\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{3})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ (105)
@@ IV Conclusions
Furthermore, we have extended this Hamiltonian to a 3D altermagnetic one preserving the $C_{4}^{z}\mathbb{T}$ symmetry. In the negligible SOC limit, our analysis reveals the presence of two Weyl points. However, when incorporating SOC, we observe the emergence of eight Weyl points and two spin Weyl points. In these cases, our model defines a Weyl altermagnetic semimetal, in which the spin Chern signal is linked to the locations of both the Weyl and spin Weyl points. The presence of these Weyl points confirms that the material exhibits non-trivial topological features characteristic of a Weyl semimetal, opening pathways to explore exotic phenomena such as the chiral anomaly and surface state physics. Finally, the formulation of a $C_{4}^{z}\mathbb{T}$ topological invariant presented in this work can be extended to other altermagnetic systems, where spin-opposite sublattices are connected through a combination of rotation and time-reversal symmetry.
@@ Acknowledgments
RGH gratefully acknowledges the computing time granted on the supercomputer Mogon at Johannes Gutenberg University Mainz (hpc.uni-mainz.de). Additionally, the support from the Universidad Nacional de Colombia (QUIPU code 202010042199) and from MinCiencias through Convocatoria 937 for Fundamental Research is also deeply appreciated. HS acknowledges the support of CONAHCyT through grant number CVU 926934. BU acknowledges the support of Max Planck Institute for Mathematics in Bonn, Germany. RGH and BU thank the continuous support of the Alexander Von Humboldt Foundation, Germany.